The length of a rectangular frame is 8cm more than the width. The area inside the frame is 105 sq cm. Find the width of the frame.
Answer by checkley71(8403)

105=2*(W+8)+2W OR 105=2W+16+2W OR 105=4W+16 OR 4W=89 OR W=89/4 OR W=22.25